OpenStudy (anonymous):

4x+5 over 3 + 2x-1 over 5 = 2 solve

OpenStudy (anonymous):

\[\frac{4x+5}{3}+\frac{2x-1}{5}=2\] LCD is 15: \[\frac{5(4x+5)}{15}+\frac{3(2x-1)}{15}=2\] \[\frac{20x+25+6x-3}{15}=2\] \[\frac{26x+22}{15}=2\] \[26x+22=30\] \[26x=8\] \[x=\frac{4}{13}\]